Assume that MKU Enterprise Ltd sells a 500 KES bond that matures in 10 years. The annual interest payments are 45 KES and the net proceeds of the bond are Ksh. 478. Use the correct formula and the above given data to estimate before-tax-cost of bond.  


1
Expert's answer
2022-02-13T11:48:16-0500

"r=\frac{\frac{500-478}{10}+45}{\frac{500+478}{2}}=\frac{47.2}{489}=0.0965" or 9.65%